OpenStudy (gazela1997):

ashur traveled 100 km/h from oakville to oshawa by car and go train. The car averaged 60km/h and the train averaged 90km/h. The whole trip took 1.5 hours. How much time did he spent on the train. HELP MEDAL AND FAN

OpenStudy (gazela1997):

btw this is grade 10 math linear systems

OpenStudy (gazela1997):

so I need 2 linear equations

OpenStudy (gazela1997):

@satellite73

OpenStudy (gazela1997):

@superhelp101

OpenStudy (superhelp101):

ugh I have no clue i am only in 9th grade :/ @iPwnBunnies do you know how to do this?

OpenStudy (superhelp101):

@matt101

OpenStudy (gazela1997):

I need 2 linear equations to solve this

OpenStudy (matt101):

Yup I know...you get them by making that chart. I just had a mistake in it - gimme a sec

OpenStudy (gazela1997):

ok

OpenStudy (gazela1997):

can you type the equations?The only problem i have is i don't know how to write equations for this

OpenStudy (matt101):

You'll see where I'm going with this in a sec :P It's worth it I promise! Also I just tweaked the chart a bit to make it easier to solve, so look at this one instead:

OpenStudy (gazela1997):

ok

OpenStudy (matt101):

I took the information from the question and laid it out in a way that's a bit easier to understand. The trip has two parts - a part traveled by car and a part traveled by train. There are different distances, speeds, and times involved with each one, so I separated them. We know the total time traveled (1.5 hours), but we want to find the time traveled by train only. I called the time traveled by car x and the time traveled by train y. This gives us our first equation: x+y=1.5 The question tells us the total distance traveled (I think you meant 100 km, not 100 km/h), but what we don't know are the individual distances traveled by car and by train. But remember S=D/T? If we know any two of S, D, or T, we can calculate the third. That's how I got the values in the first column: S x T = D. This gives us our second equation: 60x+90y=100 I think you can work it from there. You can see the equations come from just reading down the columns in the chart. Whenever you have these kinds of questions, make a chart, and the equations will pop out! Sorry it took so long to get this out lol...